OpenStudy (anonymous):

Find the surface area of each right prism with the given features. Round to the nearest whole number. a. The bases are equilateral triangles with sides of length 8 ft, and the height of the prism is 10 ft. b. The bases are trapezoids with parallel sides of lengths 7 cm and 9 cm perpendicular to one side of length 6 cm. The height of the prism is 12 cm.

OpenStudy (s):

 So you have 2 triangles and 3 rectangles.. find the area of each and add up.. (area of a triangle) times 2 (cause you have 2 of them) + (area of a rectangle) times 3 (cause you have 3 of them)  a^2 + b^2 = c^2 4^2 + h^2 = 8^2 16 + h^2 = 64 h^2 = 48 h = sqrt(48) Now we can find the area of the triangle (1/2)(base)(height) = (1/2)*(4)*(sqrt(48)) = 13.86 Now lets find the area of a rectangle Side * side = 8 * 10 = 80 Now lets plug in to our original formula 2*(13.86) + 3*(80) = 267.72 = 268 square feet